OpenStudy (anonymous):

I got most part of the problem..but not yet all of it.. :( it'd be nice if osmeone could help me with this agian..

OpenStudy (anonymous):

(2i)^2 = 2n (n+1)(2n+1)/3 a)Show true for n=1 4 = 4 b) Show true for n = k 4 + 16 +36 + 64 + (2k)^2 = 2k (k+1)(2k+1)/3 c) 4 + 16 +36 + 64 + (2k)^2 + (2(k+1))^2 = 2(k+1)((k+1)+1)(2(k+1)+1)/3 2k (k+1)(2k +1)/3 + (2(k+1))^2 = 2(k+1)((k+1)+1)(2(k+1)+1)/3 2k (k+1)(2k+1)/3 + (2(k+1))^2 (3)/3 = I got this at the end: 4k^3 + 18k^2 +26k + 12/3 (which is correct.. but I dont know where to go from here..)

OpenStudy (anonymous):

2i^2 ?

OpenStudy (anonymous):

(2i)^2

OpenStudy (anonymous):

Is always -4

OpenStudy (anonymous):

i stands for n. not imaginary

OpenStudy (anonymous):

OK, (2n)^2 and you want an induction proof, right?

OpenStudy (anonymous):

um yeah..like the way i showed it..kidna..but i got stuck at the end.. :(

OpenStudy (anonymous):

\[\begin{align} \dfrac{2k (k+1)(2k +1)}{3} + (2(k+1))^2 &= \dfrac{2(k+1)((k+1)+1)(2(k+1)+1)}{3}\\ \dfrac{(2k^2+2k)(2k +1)}{3} + (2k+2)^2 &= \dfrac{2(k+1)(k+2)(2k+3)}{3}\\ \dfrac{4k^3+6k^2+2k}{3} + \dfrac{3(2k+2)^2}{3} &= \dfrac{(2k^2+6k+4)(2k+3)}{3}\\ \dfrac{4k^3+6k^2+2k}{3} + \dfrac{3(4k^2+8k+4)}{3} &= \dfrac{4k^3+18k^2+26k+12}{3}\\ \dfrac{4k^3+6k^2+2k}{3} + \dfrac{12k^2+24k+12}{3} &= \dfrac{4k^3+18k^2+26k+12}{3}\\ \dfrac{4k^3+18k^2+26k+12}{3} &= \dfrac{4k^3+18k^2+26k+12}{3} \end{align}\]

OpenStudy (anonymous):

whoaa.

OpenStudy (anonymous):

Assume true for k, does that imply true for k+1

OpenStudy (anonymous):

Oh..nbouscal..we'arent allowd to touch the right side..

OpenStudy (anonymous):

Yes it does..

OpenStudy (anonymous):

Then you are done.

OpenStudy (anonymous):

um...u have to prove it alegbraically..

OpenStudy (anonymous):

Bah. You can always touch the right side. Anyway, just do the inverse of anything I did to the right side to the left side instead and then you're good.

OpenStudy (anonymous):

okay lemme try.. (lol its my teachers pet peeve on test to only mess with the left side and not the rite.)

OpenStudy (anonymous):

Your teacher's an idiot, then. (in my humble opinion)

OpenStudy (anonymous):

*Best Answer*^

OpenStudy (anonymous):

Im sorry if i m bothering you..but i tried it backwards, and tried doing synthetic division, firstdividng it by k+1 to get 4k^2 +14k + 12 and divind that by k+2, to get4k +6.. which: (k+1)(k+2)(4k+6) (k+1) ((k+1)+1)(2(2k+3)/3..... and not what i needed to get.. :( sorry!1 :(

OpenStudy (anonymous):

\[ \begin{align} \dfrac{2k (k+1)(2k +1)}{3} + (2(k+1))^2 &= \dfrac{2(k+1)((k+1)+1)(2(k+1)+1)}{3}\\ \dfrac{(2k^2+2k)(2k +1)}{3} + (2k+2)^2 &= \\ \dfrac{4k^3+6k^2+2k}{3} + \dfrac{3(2k+2)^2}{3} &=\\ \dfrac{4k^3+6k^2+2k}{3} + \dfrac{3(4k^2+8k+4)}{3} &= \\ \dfrac{4k^3+6k^2+2k}{3} + \dfrac{12k^2+24k+12}{3} &= \\ \dfrac{4k^3+18k^2+26k+12}{3} &=\\ \dfrac{(2k^2+6k+4)(2k+3)}{3}&=\\ \dfrac{2(k+1)(k+2)(2k+3)}{3}&=\\ \dfrac{2(k+1)((k+1)+1)(2(k+1)+1)}{3}&=\dfrac{2(k+1)((k+1)+1)(2(k+1)+1)}{3}\\ \end{align} \]

OpenStudy (anonymous):

Um...one last question.. : How'd you get 2k^2 + 6k + 4 and 2k+3?